# XMAD-Bench: Cross-Domain Multilingual Audio Deepfake Benchmark

### v1

@@ 3. Dataset
Finally, all real and fake clips are trimmed for silence at the start and the end, and resampled to 16 kHz. This ensures that real and fake samples cannot be distinguished based on the sampling rate or the presence of silence periods.
Statistics.
To assess the quality differences between real and fake samples, we report the speech activity ratio (SAR) and the signal-to-noise ratio (SNR) in Table 2. We highlight that the SAR and SNR values are fairly similar for both real and fake samples, confirming that the fake samples are of high quality.
4. Experiments
Audio deepfake detectors.
For deepfake detection, we consider four pre-trained models, namely ResNet-18 and ResNet-50 (He et al., 2016) from Torchvision444https://github.com/pytorch/vision, Audio Spectrogram Transformer (AST) (Gong et al., 2021) from Hugging Face555https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593, and wav2vec 2.0 (Baevski et al., 2020; Tak et al., 2022) from Hugging Face666https://huggingface.co/docs/transformers/model_doc/wav2vec2. We fine-tune these models on XMAD-Bench. We also train a fifth model from scratch, namely SepTr (Ristea et al., 2022). The SepTr model777https://github.com/ristea/septr is composed of 3 separable transformer blocks, each with 5 attention heads. The dimension of each head and the dimension of the MLP layer are set to 256. Each audio clip is converted into a spectrogram to be processed by each of the five models.
Experimental setup.
We conduct experiments by training detectors on the training set of each language, and evaluating them on the in-domain test set, after every epoch. The checkpoints achieving the highest performance on the in-domain split are further tested on the cross-domain test set. In addition, we also carry out cross-lingual experiments, training the models on Arabic, German, Romanian, Russian and Spanish, and testing them on English and Mandarin. In the cross-lingual setup, we randomly select at most 3,000 samples per language. In all experiments, each clip is augmented during training with a probability of . Augmentations include time shifting by rolling the signal, speed augmentation, volume augmentation by applying random gain, clipping, reverberation, spectral shifting (high-shelf, low-shelf and peak filters) and pitch shifting.
Hyperparameters.
@@ 4. Experiments
A figure showing the cross-domain confusion matrices of ResNet-18 (first row) and AST (second row) on Arabic (first column) and English (second column).
Results.
As shown in Table 3, several models reach an impressive performance of 100% on the in-domain split, confirming that deep neural models can easily capture the characteristics of generative models, regardless of the target language. However, the cross-domain results support our conjecture, demonstrating that most metrics drastically decrease when detectors are tested in the cross-domain scenario, where fake samples are generated by models unknown to the detectors. Notably, wav2vec 2.0 outperforms all other models in the cross-domain setting on four of the seven target languages, showing its strong ability to generalize to unseen generative models. Furthermore, wav2vec 2.0 exhibits robust multilingual performance, ranking among the top performers in the multilingual setting and demonstrating its effectiveness in capturing language-independent representations of synthetic speech. The cross-lingual experiments indicate that the language shift does not significantly affect performance.
@@ Results.
5. Conclusion and Future Work
In this paper, we introduced a novel multilingual cross-domain audio dataset to evaluate deepfake detectors “in the wild”. We discussed the methodology used to generate diverse fake clips for three partitions (training, in-domain test, and cross-domain test), and we further presented dataset statistics across the seven target languages. We evaluated the performance of five state-of-the-art models from the literature in terms of multiple metrics, showing that our cross-domain evaluation scenario causes a significant decline for all the reported metrics. While we were able to replicate the near perfect in-domain performance previously reported in literature, our cross-domain results highlighted the difficulty of performing audio deepfake detection across datasets, languages, speakers, and deepfake generative methods.
@@ Acknowledgements.
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-2-0227, within PNCDI IV.References
- (1)
- Al-Fetyani et al. (2023) Mohammad Al-Fetyani, Muhammad Al-Barham, Gheith Abandah, Adham Alsharkawi, and Maha Dawas. 2023. MASC: Massive Arabic Speech Corpus. In SLT. 1006–1013.
- Ardila et al. (2020) R. Ardila, M. Branson, K. Davis, M. Henretty, M. Kohler, J. Meyer, R. Morais, L. Saunders, F. M. Tyers, and G. Weber. 2020. Common Voice: A Massively-Multilingual Speech Corpus. In LREC. 4218–4222.
- Baas et al. (2023) Matthew Baas, Benjamin van Niekerk, and Herman Kamper. 2023. Voice Conversion With Just Nearest Neighbors. In INTERSPEECH. 2053–2057.
- Baevski et al. (2020) Alexei Baevski, Yuhao Zhou, Abdelrahman Mohamed, and Michael Auli. 2020. wav2vec 2.0: A framework for self-supervised learning of speech representations. In NeurIPS, Vol. 33. 12449–12460.
- Bu et al. (2017) Hui Bu, Jiayu Du, Xingyu Na, Bengu Wu, and Hao Zheng. 2017. AISHELL-1: An Open-Source Mandarin Speech Corpus and A Speech Recognition Baseline. In O-COCOSDA.
- Casanova et al. (2024) Edresson Casanova, Kelly Davis, Eren Gölge, Görkem Göknar, Iulian Gulea, Logan Hart, Aya Aljafari, Joshua Meyer, Reuben Morais, Samuel Olayemi, and Julian Weber. 2024. XTTS: a Massively Multilingual Zero-Shot Text-to-Speech Model. In INTERSPEECH. 4978–4982.
- Casanova et al. (2022) Edresson Casanova, Julian Weber, Christopher D. Shulby, Arnaldo Candido Junior, Eren Gölge, and Moacir A. Ponti. 2022. YourTTS: Towards Zero-Shot Multi-Speaker TTS and Zero-Shot Voice Conversion for Everyone. In ICML. 2709–2720.
- Chen et al. (2023) Feng Chen, Shiwen Deng, Tieran Zheng, Yongjun He, and Jiqing Han. 2023. Graph-based spectro-temporal dependency modeling for anti-spoofing. In ICASSP.
- Chen et al. (2025) Sanyuan Chen, Chengyi Wang, Yu Wu, Ziqiang Zhang, Long Zhou, Shujie Liu, Zhuo Chen, Yanqing Liu, Huaming Wang, Jinyu Li, Lei He, Sheng Zhao, and Furu Wei. 2025. Neural Codec Language Models are Zero-Shot Text to Speech Synthesizers. IEEE Transactions on Audio, Speech and Language Processing 33 (2025), 705–718.
- Croitoru et al. (2024) Florinel-Alin Croitoru, Andrei-Iulian Hiji, Vlad Hondru, Nicolae Catalin Ristea, Paul Irofti, Marius Popescu, Cristian Rusu, Radu Tudor Ionescu, Fahad Shahbaz Khan, and Mubarak Shah. 2024. Deepfake Media Generation and Detection in the Generative AI Era: A Survey and Outlook. arXiv preprint arXiv:2411.19537 (2024).
- Frank and Schönherr (2021) J. Frank and L Schönherr. 2021. WaveFake: A Data Set to Facilitate Audio Deepfake Detection. In NeurIPS.
- Fu et al. (2021) Yihui Fu, Luyao Cheng, Shubo Lv, Yukai Jv, Yuxiang Kong, Zhuo Chen, Yanxin Hu, Lei Xie, Jian Wu, Hui Bu, et al. 2021. AISHELL-4: An Open Source Dataset for Speech Enhancement, Separation, Recognition and Speaker Diarization in Conference Scenario. In INTERSPEECH. 3665–3669.
- Gong et al. (2021) Yuan Gong, Yu-An Chung, and James Glass. 2021. AST: Audio Spectrogram Transformer. In INTERSPEECH. 571–575.
- He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep Residual Learning for Image Recognition. In CVPR. 770–778.
- Huang et al. (2022) R. Huang, M.W.Y. Lam, J. Wang, D. Su, D. Yu, Y. Ren, and Z. Zhao. 2022. FastDiff: A fast conditional diffusion model for high-quality speech synthesis. In IJCAI. 4157–4163.
- Hung et al. (2022) Yun-Ning Hung, Chih-Wei Wu, Iroro Orife, Aaron Hipple, William Wolcott, and Alexander Lerch. 2022. A Large TV Dataset for Speech and Music Activity Detection. EURASIP Journal on Audio, Speech, and Music Processing 2022, 1 (2022), 21.
- Ito and Johnson (2017) Keith Ito and Linda Johnson. 2017. The LJ Speech Dataset. https://keithito.com/LJ-Speech-Dataset/.
- Jiang et al. (2023) Ziyue Jiang, Yi Ren, Zhenhui Ye, Jinglin Liu, Chen Zhang, Qian Yang, Shengpeng Ji, Rongjie Huang, Chunfeng Wang, Xiang Yin, et al. 2023. Mega-TTS: Zero-Shot Text-to-Speech at Scale with Intrinsic Inductive Bias. arXiv preprint arXiv:2306.03509 (2023).
- Ju et al. (2024) Zeqian Ju, Yuancheng Wang, Kai Shen, Xu Tan, Detai Xin, Dongchao Yang, Eric Liu, Yichong Leng, Kaitao Song, Siliang Tang, et al. 2024. NaturalSpeech 3: Zero-Shot Speech Synthesis with Factorized Codec and Diffusion Models. In ICML.
- Jung et al. (2022) Jee-weon Jung, Hee-Soo Heo, Hemlata Tak, Hye-jin Shim, Joon Son Chung, Bong-Jin Lee, Ha-Jin Yu, and Nicholas Evans. 2022. AASIST: Audio Anti-Spoofing using Integrated Spectro-Temporal Graph Attention Networks. In ICASSP. 6367–6371.
- Kim et al. (2020) Jaehyeon Kim, Sungwon Kim, Jungil Kong, and Sungroh Yoon. 2020. Glow-tts: A generative flow for text-to-speech via monotonic alignment search. Advances in Neural Information Processing Systems 33 (2020), 8067–8077.
- Kim et al. (2021) Jaehyeon Kim, Jungil Kong, and Juhee Son. 2021. Conditional Variational Autoencoder with Adversarial Learning for End-to-End Text-to-Speech. In ICML. 5530–5540.
- Lee et al. (2023) Sang-gil Lee, Wei Ping, Boris Ginsburg, Bryan Catanzaro, and Sungroh Yoon. 2023. BigVGAN: A Universal Neural Vocoder with Large-Scale Training. In ICLR.
- Li et al. (2022) Jingyi Li, Weiping Tu, and Li Xiao. 2022. FreeVC: Towards High-Quality Text-Free One-Shot Voice Conversion. In ICASSP.
- Li et al. (2024) Yuang Li, Min Zhang, Mengxin Ren, Xiaosong Qiao, Miaomiao Ma, Daimeng Wei, and Hao Yang. 2024. Cross-Domain Audio Deepfake Detection: Dataset and Analysis. In EMNLP. 4977–4983.
- Liu et al. (2023) Xiaohui Liu, Meng Liu, Longbiao Wang, Kong Aik Lee, Hanyi Zhang, and Jianwu Dang. 2023. Leveraging positional-related local-global dependency for synthetic speech detection. In ICASSP.
- Müller et al. (2022) Nicolas M Müller, Pavel Czempin, Franziska Dieckmann, Adam Froghyar, and Konstantin Böttinger. 2022. Does audio deepfake detection generalize?. In INTERSPEECH. 2783–2787.
- Müller et al. (2024) Nicolas M Müller, Piotr Kawa, Wei Herng Choong, Edresson Casanova, Eren Gölge, Thorsten Müller, Piotr Syga, Philip Sperl, and Konstantin Böttinger. 2024. MLAAD: The Multi-Language Audio Anti-Spoofing Dataset. In IJCNN.
- Pratap et al. (2024) Vineel Pratap, Andros Tjandra, Bowen Shi, Paden Tomasello, Arun Babu, Sayani Kundu, Ali Elkahky, Zhaoheng Ni, Apoorv Vyas, Maryam Fazel-Zarandi, Alexei Baevski, Yossi Adi, Xiaohui Zhang, Wei-Ning Hsu, Alexis Conneau, and Michael Auli. 2024. Scaling Speech Technology to 1,000+ Languages. Journal of Machine Learning Research 25, 97 (2024), 1–52.
- Qin et al. (2024) Zengyi Qin, Wenliang Zhao, Xumin Yu, and Xin Sun. 2024. OpenVoice: Versatile Instant Voice Cloning. arXiv preprint arXiv:2312.01479 (2024).
- Radford et al. (2023) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever. 2023. Robust Speech Recognition via Large-Scale Weak Supervision. In ICML. 28492–28518.
- Reimao and Tzerpos (2019) Ricardo Reimao and Vassilios Tzerpos. 2019. FoR: A dataset for synthetic speech detection. In SpeD. 1–10.
- Ristea et al. (2022) Nicolae-Cătălin Ristea, Radu Tudor Ionescu, and Fahad Shahbaz Khan. 2022. SepTr: Separable Transformer for Audio Spectrogram Processing. In INTERSPEECH. 4103–4107.
- Roselló Casado et al. (2023) Eros Roselló Casado, Alejandro Gómez Alanís, Ángel Manuel Gómez García, Antonio Miguel Peinado Herreros, et al. 2023. A conformer-based classifier for variable-length utterance processing in anti-spoofing. In INTERSPEECH. 5281–5285.
- Shen et al. (2018) Jonathan Shen, Ruoming Pang, Ron J. Weiss, Mike Schuster, Navdeep Jaitly, Zongheng Yang, Zhifeng Chen, Yu Zhang, Yuxuan Wang, Rj Skerrv-Ryan, Rif A. Saurous, Yannis Agiomvrgiannakis, and Yonghui Wu. 2018. Natural TTS Synthesis by Conditioning Wavenet on MEL Spectrogram Predictions. In ICASSP. 4779–4783.
- Shen et al. (2024) Kai Shen, Zeqian Ju, Xu Tan, Yanqing Liu, Yichong Leng, Lei He, Tao Qin, Sheng Zhao, and Jiang Bian. 2024. NaturalSpeech 2: Latent Diffusion Models are Natural and Zero-Shot Speech and Singing Synthesizers. In ICLR.
- Shi et al. (2021) Yao Shi, Hui Bu, Xin Xu, Shaoji Zhang, and Ming Li. 2021. AISHELL-3: A Multi-Speaker Mandarin TTS Corpus. In INTERSPEECH. 2756–2760.
- Solak (2019) I. Solak. 2019. The M-AILABS Speech Dataset.
- Sonobe et al. (2017) Ryosuke Sonobe, Shinnosuke Takamichi, and Hiroshi Saruwatari. 2017. JSUT corpus: free large-scale Japanese speech corpus for end-to-end speech synthesis. arXiv preprint arXiv:1711.00354 (2017).
- Suno (2023) Suno. 2023. Bark. https://github.com/suno-ai/bark
- Tak et al. (2022) Hemlata Tak, Massimiliano Todisco, Xin Wang, Jee-weon Jung, Junichi Yamagishi, and Nicholas Evans. 2022. Automatic speaker verification spoofing and deepfake detection using Wav2Vec 2.0 and data augmentation. In Odyssey. 112–119.
- Tan et al. (2024) Xu Tan, Jiawei Chen, Haohe Liu, Jian Cong, Chen Zhang, Yanqing Liu, Xi Wang, Yichong Leng, Yuanhao Yi, Lei He, Sheng Zhao, Tao Qin, Frank Soong, and Tie-Yan Liu. 2024. NaturalSpeech: End-to-End Text-to-Speech Synthesis With Human-Level Quality. IEEE Transactions on Pattern Analysis and Machine Intelligence 46, 6 (2024), 4234–4245.
- Truong et al. (2024) Duc-Tuan Truong, Ruijie Tao, Tuan Nguyen, Hieu-Thi Luong, Kong Aik Lee, and Eng Siong Chng. 2024. Temporal-Channel Modeling in Multi-head Self-Attention for Synthetic Speech Detection. In INTERSPEECH. 537–541.
- Wang et al. (2021) Changhan Wang, Morgane Riviere, Ann Lee, Anne Wu, Chaitanya Talnikar, Daniel Haziza, Mary Williamson, Juan Pino, and Emmanuel Dupoux. 2021. VoxPopuli: A Large-Scale Multilingual Speech Corpus for Representation Learning, Semi-Supervised Learning and Interpretation. In ACL. 993–1003.
- Wang et al. (2020) Xin Wang, Junichi Yamagishi, Massimiliano Todisco, Héctor Delgado, Andreas Nautsch, Nicholas Evans, Md Sahidullah, Ville Vestman, Tomi Kinnunen, Kong Aik Lee, et al. 2020. ASVspoof 2019: A large-scale public database of synthesized, converted and replayed speech. Computer Speech & Language 64 (2020), 101114.
- Yamagishi et al. (2017) Junichi Yamagishi, Christophe Veaux, and Kirsten MacDonald. 2017. CSTR VCTK Corpus: English Multi-speaker Corpus for CSTR Voice Cloning Toolkit. https://doi.org/10.7488/ds/1994 University of Edinburgh. The Centre for Speech Technology Research (CSTR).
- Yamagishi et al. (2021) Junichi Yamagishi, Xin Wang, Massimiliano Todisco, Md Sahidullah, Jose Patino, Andreas Nautsch, Xuechen Liu, Kong Aik Lee, Tomi Kinnunen, Nicholas Evans, et al. 2021. ASVspoof 2021: accelerating progress in spoofed and deepfake speech detection. In ASVspoof. 47–54.
- Yi et al. (2022) Jiangyan Yi, Ruibo Fu, Jianhua Tao, Shuai Nie, Haoxin Ma, Chenglong Wang, Tao Wang, Zhengkun Tian, Ye Bai, Cunhang Fan, et al. 2022. ADD 2022: The First Audio Deep Synthesis Detection Challenge. In ICASSP. 9216–9220.
- Yi et al. (2023) Jiangyan Yi, Jianhua Tao, Ruibo Fu, Xinrui Yan, Chenglong Wang, Tao Wang, Chu Yuan Zhang, Xiaohui Zhang, Yan Zhao, Yong Ren, et al. 2023. ADD 2023: The Second Audio Deepfake Detection Challenge. arXiv preprint arXiv:2305.13774 (2023).
- Zhang et al. (2023) Ziqiang Zhang, Long Zhou, Chengyi Wang, Sanyuan Chen, Yu Wu, Shujie Liu, Zhuo Chen, Yanqing Liu, Huaming Wang, Jinyu Li, et al. 2023. Speak foreign languages with your own voice: Cross-lingual neural codec language modeling. arXiv preprint arXiv:2303.03926 (2023).
- Zhao et al. (2023) Wenliang Zhao, Xumin Yu, and Zengyi Qin. 2023. MeloTTS: High-quality Multi-lingual Multi-accent Text-to-Speech. https://github.com/myshell-ai/MeloTTS

### v2

@@ 4.2 Experimental Setup
| Whisper+MLP | 99.75 ( 0.02) | 99.99 ( 0.00) | 0.24 ( 0.01) | 72.80 ( 0.67) | 90.22 ( 0.79) | 17.77 ( 1.10) | |
4.3 Hyperparameters
@@
4.4 Main Results
As shown in Table 3, several models reach an impressive performance of 100% on the in-domain split, confirming that deep neural models can easily capture the characteristics of generative models, regardless of the target language. However, the cross-domain results support our conjecture, demonstrating that most metrics drastically decrease when detectors are tested in the cross-domain scenario, where fake samples are generated by models unknown to the detectors. Notably, wav2vec 2.0 and Whishper+MLP outperform all other models in the cross-domain setting on most target languages, both showing a strong ability to generalize to unseen generative models. Furthermore, wav2vec 2.0 exhibits robust multilingual performance, ranking among the top performers in the multilingual setting and demonstrating its effectiveness in capturing language-independent representations of synthetic speech. The cross-lingual in-domain experiments indicate that the language shift does not significantly affect performance. The performance still degrades in the cross-lingual cross-domain scenario.
@@ 4.4 Main Results
4.5 Results for Background Noise Injection
We further evaluate how one of the best performing models handles the deepfake detection task under background noise injection. More specifically, we employ the Whisper-based model on Romanian test samples augmented with Gaussian noise, using two values for the standard deviation of the added noise, namely and . Both noise levels can be heard by humans, but the second one makes it difficult to distinguish some words.
In Table 4, we show the cross-domain results of the Whisper+MLP model trained on clean samples (without noise injection). The results indicate that the introduction of audible yet moderate noise () does not affect performance. Increasing the noise magnitude () degrades performance, but at this noise level, some utterances are indistinguishable by humans. Overall, we conclude that Whisper+MLP obtains reasonable performance when samples are affected by noise, likely because the pre-trained Whisper-Large-v3 backbone is robust to background noise injection.
| Noise level | ACC | AUC | EER |
| (no noise) | 75.72 | 99.54 | 2.17 |
| 80.77 | 98.78 | 4.64 | |
| 68.70 | 98.29 | 5.93 |
5 Conclusion and Future Work
In this paper, we introduced a novel multilingual cross-domain audio dataset to evaluate deepfake detectors “in the wild”. We discussed the methodology used to generate diverse fake clips for three partitions (training, in-domain test, and cross-domain test), and we further presented dataset statistics across the seven target languages. We evaluated the performance of six state-of-the-art models from the literature in terms of multiple metrics, showing that our cross-domain evaluation scenario causes a significant decline for all the reported metrics. While we were able to replicate the near perfect in-domain performance previously reported in literature, our cross-domain results highlighted the difficulty of performing audio deepfake detection across datasets, languages, speakers, and deepfake generative methods.
@@ 6 Acknowledgments
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-2-0227, within PNCDI IV. This research is also supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 351416.
7 Limitations
To construct our audio deepfake detection benchmark, we relied on recent and publicly available text-to-speech and voice conversion methods. Unfortunately, most existing TTS and VC methods do not offer support for all the targeted languages. Hence, we were forced to use distinct generative methods across the chosen languages. Nevertheless, we kept in-domain vs. cross-domain separation across all languages.
Our benchmark provides a challenging cross-domain evaluation setup, where noticeable performance drops can be observed. Yet, we did not try to adapt deepfake detectors to the cross-domain evaluation setting. Adapting models for the cross-domain setting is a challenge that requires careful consideration. We believe this exploration is beyond the goal of constructing a challenging benchmark for audio deepfake detection, so we leave it for future research.
8 Potential Risks
The development of audio deepfake generation models can have significant implications for our society, as it facilitates the spread of misinformation and phishing attacks. As synthetic audio becomes increasingly realistic and accessible, the risk of misuse continues to grow. To fight against this, more competent detection models are required. Challenging dataset construction represents one way to advance research on robust detection systems, as such models heavily depend on the utilized training data. Our benchmark fosters the development of audio deepfake detectors, as it addresses some of the limitations of previous datasets: a variety of generation methods and languages, as well as a meticulously designed cross-domain split. Nevertheless, we acknowledge that the development of detection methods can inadvertently push research towards more sophisticated generative models.
9 License
We share XMAD-Bench under the International Attribution Non-Commercial Share-Alike 4.0 (CC BY-NC-SA 4.0)888https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en license, aiming for open and responsible research on deepfake detection.